OpenStudy (wintersuntime):

Solving this problem using elimination x+5y=8 -x +2y= -1

OpenStudy (wintersuntime):

@mathmale

OpenStudy (aveline):

You can solve this equation like an addition problem. Solving vertically, eliminating the x's first (one is positive and one is negative so they cancel out) x+5y=8 -x+2y=-1 ------------- 7y=7 Simplify: y=1 Now, you can just substitute 1 for y: x+5=8 x=3

OpenStudy (wintersuntime):

thanks